# Towards a Science of Scaling Agent Systems

### v1

@@ Abstract
Figure 1: Agent Scaling across model intelligence and system architectures. Average performance (%) across four agentic benchmarks improves consistently with increasing model Intelligence Index (see Appendix A) across three major LLM families (OpenAI, Google, and Anthropic) under different agent configurations. Single Agent System (SAS) serves as reference trajectories, while Multi Agent System (MAS) variants (Centralized, Decentralized, Independent, and Hybrid) reveal distinct scaling behaviors (see Table 2 for architecture comparisons). All percentage deltas annotated in the figure (e.g., +8.1%, +8.7%, –4.6%) indicate relative performance change of the best-performing MAS variant compared to the SAS baseline at the same Intelligence Index. Centralized and hybrid coordination generally yield superior scaling efficiency, suggesting that collaborative agentic structures amplify capability gains more effectively than individual scaling alone.
@@ 1 Introduction
Agents (wang2024survey), language model-driven systems that operate through iterative cycles of reasoning, planning, and acting, adapting their behavior based on environmental or tool-generated feedback, have achieved remarkable performance in diverse applications, from code generation (zhang2024codeagent; yang2024swe), web browsing (wei2025browsecomp; yao2022webshop), medical decision-making (heydari2025anatomy; mcduff2025towards; kim2024mdagents), finance (yu2025finmem), to scientific discovery (gottweis2025towards; mitchener2025kosmos). As tasks grow in complexity and require sustained environmental interaction, the field has increasingly turned to multi-agent systems (MAS), relying on the premise that specialized collaboration consistently outperforms single-agent systems (SAS) (tran2025multiagent; guo2024large). Previous work has made positive claims about multi-agent systems: “More agents is all you need(li2024more), suggesting that agent collaboration follows collaborative scaling principles (qian2024scaling), and that MAS consistently outperforms single-agent systems (SAS) on complex tasks (du2023improving; chen2023agentverse). Yet, despite rapid adoption, there remains no principled quantitative framework to predict when adding agents amplifies performance and when it erodes it. This gap leaves practitioners relying on heuristics, hindering both the emergence of a science of agent systems and, critically for real-world deployment, the ability to determine when multi-agent coordination provides genuine value over simpler single-agent alternatives.
@@
In contrast to prior claims that “more agents is all you need”, our evaluation reveals that the effectiveness of multi-agent systems is governed by quantifiable trade-offs between architectural properties and task characteristics. We establish a predictive framework using empirical coordination metrics: efficiency (success/overhead ratio), error amplification factors, message density and redundancy, achieving cross-validated $R^{2}{=}0.513$ (explaining more than half of the performance variance on held-out data) without dataset-specific parameters. Critically, this framework generalizes beyond training configurations: leave-one-domain-out cross-validation achieves $R^{2}{=}0.89$, and the model correctly predicts optimal architectures for 87% of held-out task configurations, demonstrating extrapolation to unseen task structures.
@@
Our analysis identifies three patterns. First, a tool-coordination trade-off ($\beta{=}{-}0.330$, p<0.001p{<}0.001): tool-heavy tasks (e.g., 16-tool software engineering) suffer from multi-agent coordination overhead, with efficiency penalties compounding as environmental complexity increases. Second, a capability ceiling ($\beta{=}{-}0.408$, p<0.001p{<}0.001): tasks where single-agent performance already exceeds 45% accuracy experience negative returns from additional agents, as coordination costs exceed diminishing improvement potential. Third, architecture-dependent error amplification where independent multi-agent systems amplify errors 17.2-fold versus single-agent baseline through unchecked error propagation, where errors made by individual agents propagate to the final output without inter-agent verification, while centralized coordination achieves 4.4-fold containment via validation bottlenecks, where the orchestrator reviews sub-agent outputs before aggregation., catching errors before they propagate to the final response. Performance spans $81\%$ relative improvement (structured financial reasoning under centralized coordination) to $70\%$ degradation (sequential planning under independent coordination), demonstrating that architecture-task alignment, not number of agents, determines collaborative success. Importantly, optimal architectures vary systematically: decentralized coordination benefits tasks requiring parallel exploration of high-entropy search spaces (dynamic web navigation: $9.2\%$), while all multi-agent variants universally degrade performance on tasks requiring sequential constraint satisfaction (planning: $39\%$ to $70\%$), where coordination overhead fragments reasoning capacity under fixed computational budgets. We synthesize these findings into quantitative architecture selection rules (Section 4.3) achieving 87% prediction accuracy on held-out configurations. The underlying mechanisms driving these patterns are interpretable: the tool-coordination trade-off arises because multi-agent systems fragment the per-agent token budget, leaving insufficient capacity for complex tool orchestration; the capability ceiling reflects that coordination overhead becomes a net cost when baseline performance is already high; and architecture-dependent error amplification stems from the presence or absence of validation bottlenecks that catch errors before propagation. These mechanistic insights enable practitioners to move from architectural heuristics to principled, measurement-driven deployment decisions.
@@
  • Scaling principle for multi-agent systems: We derive a mixed-effects model ($R^{2}{=}0.513$) using empirical coordination metrics, including efficiency ($E_{c}$), error amplification ($A_{e}$), and redundancy ($\rho$) to quantify how performance emerges from the interplay of reasoning capability and task properties. The analysis identifies dominant inhibitory mechanisms, specifically the tool-coordination trade-off ($\beta{=}{-}0.330$) and architecture-dependent error cascades, establishing fundamental limits on coordination effectiveness.
  • Quantitative principles for architecture-task alignment: We demonstrate that agent architecture selection is governed by measurable task features (e.g., decomposability, tool complexity) rather than simple agent scaling. Our framework achieves 87% accuracy in predicting optimal architectures on held-out tasks, enabling principled deployment decisions grounded in predictive models rather than qualitative heuristics.
@@ 3.1 System Definition
The observation history space $\mathcal{H}$ contains sequences of action-observation pairs. The decision function $\pi_{i}$ is instantiated by the reasoning policy $\Phi_{i}$ (the LLM): given a history $h_{i,t}$, the LLM generates a reasoning trace and selects the next action. For instance, a history $h_{i,t}=[(\text{``search(query='pandas')''},\text{``Found 5 files''}),$...] is processed by $\Phi_{i}$ to produce the next tool call $\alpha_{i,t+1}$.
@@
  • Independent: $C=\emptyset$ (no inter-agent communication)
@@
  • Independent MAS: $A=\{a_{1},\ldots,a_{n}\}$, $C=\emptyset$, $\Omega=\text{synthesis\_only}$. Each of $n$ agents performs $k$ reasoning iterations independently, then outputs are aggregated ($O(nk)$). This achieves maximal parallelization but minimal coordination, suitable for ensemble-style reasoning.
@@ 4 Experiments & Results
Figure 2: Comparative performance of single-agent (SAS) and multi-agent systems (MAS) across four diverse benchmarks reveals highly task-dependent scaling dynamics. Box plots show distribution of success rates (scale: 0 to 1, where 1 represents 100% success). Percentage annotations represent relative improvement/degradation compared to SAS baseline: $(\text{mean}_{\text{MAS}}-\text{mean}_{\text{SAS}})/\text{mean}_{\text{SAS}}\times 100\%$. SAS serves as the reference baseline (shown without percentage annotation). (a) BrowseComp-Plus shows polarized results, with independent agents catastrophically underperforming relative to SAS (-35%) while more structured coordination achieves modest gains. (b) Finance Agent demonstrates the strongest multi-agent benefits, with all MAS architectures substantially outperforming SAS (from +57 to 81%), suggesting that complex planning and distributed reasoning provide significant advantages in structured economic domains. (c) PlanCraft exhibits consistent degradation across all MAS variants (from -70% to -39%). The core difference from Finance Agent lies in task structure, where Finance Agent tasks decompose into parallelizable subtasks (e.g., separate agents can independently analyze revenue trends, cost structures, and market comparisons, then synthesize findings), whereas PlanCraft requires strictly sequential state-dependent reasoning, each crafting action modifies the inventory state that subsequent actions depend upon. (d) Workbench shows marginal effects (from -11 to +6%), suggesting balanced trade-offs between problem structure and orchestration costs. White diamond markers denote per-architecture mean performance.
@@ 4.1 Setup
We evaluate three LLM families across multiple model sizes, spanning externally standardized Intelligence Index values from 34 to 66 (a composite capability score integrating reasoning, coding, and knowledge benchmarks; see Appendix A):
@@
We tested five coordination topologies: Single-Agent System (SAS) and Multi-Agent System (MAS) topologies: Independent (no inter-agent communication), Centralized (hub-and-spoke with orchestrator), Decentralized (peer-to-peer mesh), and Hybrid (dense but curated selective connections). Coordination complexity is parameterized as $C=E/E_{max}\cdot\log n$, where $E$ is the number of directed communication edges and $n$ the team size. This formulation reflects that (i) coordination cost scales with the number of communication channels ($E$), and (ii) message routing complexity grows logarithmically with team size under hierarchical or structured topologies, analogous to routing depth in balanced tree structures where path length scales as $O(\log n)$, consistent with distributed systems theory (malone1994interdisciplinary). For teams of 3–4 agents, this yields $C\in[0.35,1.8]$: SAS has $C=0$ (no edges), Independent has $C=0.35$ (synthesis-only aggregation), and Hybrid reaches $C=1.8$ (dense peer connections plus orchestrator).
@@
Primary outcome is task success/accuracy (domain-dependent: factual correctness for Finance Agent, task completion for Workbench, goal satisfaction for PlanCraft, page synthesis accuracy for BrowseComp-Plus). Secondary metrics include: (i) factual error rate $E$ via domain-specific validators (Cohen’s $\kappa$: Finance Agent =0.91=0.91, Workbench =0.89=0.89, PlanCraft =0.87=0.87, BrowseComp-Plus =0.88=0.88; all exceeding the substantial agreement threshold of $0.80$); (ii) information gain $\Delta\mathcal{I}$ from pre- vs. post-coordination uncertainty proxies (see Eq. 2); (iii) token-overlap structure across agent rationales, labeling tokens as unique (appearing in exactly one agent), shared (two or more agents), or contradictory (semantic opposition detected when BERTScore similarity <0.3<0.3 between assertion pairs, i.e., $1-\text{BERTScore}>0.7$, following the dissimilarity threshold established by zhang2019bertscore); (iv) efficiency metrics including success per 1,000 tokens and cost-normalized performance. All metrics are normalized per reasoning turn and per token to enable cross-architecture comparison. We select coordination metrics based on two criteria: (i) direct measurability from experimental traces without requiring ground-truth labels beyond task success, and (ii) coverage of distinct aspects of coordination–performance relationships identified in prior work (cemri2025why). We excluded metrics requiring subjective human annotation (e.g., solution creativity) or those exhibiting high collinearity with included measures (e.g., total message count correlates $r>0.92$ with overhead). Variance inflation factor (VIF) analysis confirmed no severe multicollinearity among retained predictors (all VIF <5<5). Specifically:
@@ 4.2 Main Results
Multi-agent systems demonstrate highly heterogeneous performance across task domains, contingent on problem structure and architectural choices. On Finance Agent, MAS achieve substantial improvements: Centralized reaches +80.9% (mean 0.631 vs. SAS 0.349), Decentralized achieves +74.5% (0.609), and Hybrid reaches +73.2% (0.604), driven by opportunities for distributed financial reasoning across multiple agents. On Workbench, multi-agent systems show minimal gains: Decentralized achieves +5.7% (0.664 vs. SAS 0.629), while Centralized and Hybrid both slightly underperform at -1.2%. On BrowseComp-Plus, improvements remain modest: Decentralized achieves +9.2% (0.347 vs. SAS 0.318), with Centralized essentially flat at +0.2%.
Critically, PlanCraft exhibits universal performance degradation across all multi-agent architectures. Centralized declines to -50.4% (0.282 vs. SAS 0.568), Decentralized to -41.4% (0.332), Hybrid to -39.0% (0.346), and Independent to -70.0% (0.170). This degradation reveals a coordination-saturation effect: beyond a critical reasoning–communication trade-off, additional coordination no longer improves but suppresses effective reasoning, marking the onset of communication overload. Aggregating across all benchmarks and architectures, the overall mean MAS improvement is -3.5% (95% CI: [-18.6%, +25.7%]), reflecting substantial performance heterogeneity with high variance ($\sigma=45.2\%$). The performance range across MAS variants spans from $70.0\%$ (PlanCraft Independent) to $80.9\%$ (Finance Centralized), indicating that MAS do not provide universal benefits but rather domain-specific trade-offs.
@@
Mixed-effects regression confirms domain complexity (refer to Appendix B for more details) as a significant negative moderator of MAS advantage ($\hat{\beta}=-0.114$, $95\%$ CI: [0.186,0.042][-0.186,-0.042], $p=0.002$). The mechanism operates through fixed computational budgets (matched total tokens across MAS and SAS): in structured, decomposable domains (Finance Agent, moderate Workbench instances), agents complete local reasoning with residual capacity available for inter-agent communication. Here, inter-agent messages reduce variance through redundancy elimination and enable synthesis of partial solutions, producing large performance deltas (Finance: $80.9\%$). Conversely, in high-complexity sequential domains (PlanCraft), intra-agent reasoning for constraint verification and state tracking consumes most available tokens before communication can occur; subsequent inter-agent messages then compress reasoning quality and produce strong negative returns (PlanCraft: $39.0\%$ to $70.0\%$).
@@
While domain complexity broadly moderates MAS effectiveness, the architecture-domain interaction reveals non-uniform preferences even within similar complexity regimes: no single architecture dominates across all domains and vendors. Architecture effectiveness depends critically on domain structure: Finance Agent benefits most from Centralized (+80.9%) and Decentralized (+74.5%), Workbench from MAS-Decentralized (+5.7%), and BrowseComp-Plus from MAS-Decentralized (+9.2%). In degrading domains, architecture selection becomes a least-worst optimization: PlanCraft shows Hybrid as relatively best (-39.0%) compared to MAS-Centralized (-50.4%) and MAS-Independent (-70.0%).
Family-specific coordination preferences emerge within improvement-positive domains. On Finance Agent, Anthropic’s MAS-Centralized achieves +127.5% (0.636 vs. 0.280 SAS), indicating conservative but stable coordination, whereas Google’s MAS-Centralized reaches +164.3% (0.740 vs. 0.280 SAS, averaging Centralized performance), suggesting stronger attention-mechanism alignment with hierarchical message exchange; OpenAI’s MAS-Centralized achieves +71.2% (0.79 vs. 0.465 SAS). On Workbench, where multi-agent overhead is less tolerable (efficiency degrades from $E_{c}=0.466$ for SAS to $E_{c}=0.074$ for Hybrid, the largest relative drop across benchmarks), Anthropic’s best variant (MAS-Decentralized, +10.8%) remains superior to Google (+9.5%) and OpenAI (+8.6%), reflecting relative efficiency in managing coordination costs. Critically, on PlanCraft where all variants degrade, vendor preferences flatten: Anthropic shows maximum -54.5% (MAS-Hybrid 0.31 vs. SAS 0.68), Google shows -25.3% (best), and OpenAI shows -32.3%, indicating that communication mechanisms cannot overcome fundamental sequential reasoning constraints. While the precise mechanisms remain to be characterized, potential factors include differences in instruction-following fidelity, context utilization patterns, and inter-turn consistency that affect how agents interpret and respond to coordination messages. No vendor achieves universal multi-agent dominance; instead, each exhibits relative advantages in structured domains (Finance) that evaporate in sequential constraint-satisfaction domains (PlanCraft), indicating that multi-agent benefits are genuinely contingent on problem structure rather than generalizable across task types.

### v2

@@ Abstract
Figure 1: Agent Scaling across model intelligence and system architectures. Average performance (%) across four agentic benchmarks improves consistently with increasing model Intelligence Index (see Appendix A) across three major LLM families (OpenAI, Google, and Anthropic) under different agent configurations. Single Agent System (SAS) serves as reference trajectories, while Multi Agent System (MAS) variants (Centralized, Decentralized, Independent, and Hybrid) reveal distinct scaling behaviors (see Table 2 for architecture comparisons). All percentage deltas annotated in the figure (e.g., +8.7%, +8.1%, –4.6%) indicate relative performance change of the best-performing MAS variant compared to the SAS baseline at the same Intelligence Index. Centralized and hybrid coordination generally yield superior scaling efficiency, suggesting that collaborative agentic structures amplify capability gains more effectively than individual scaling alone.
@@ 1 Introduction
Agents (wang2024survey), language model-driven systems that operate through iterative cycles of reasoning, planning, and acting, adapting their behavior based on environmental or tool-generated feedback, have achieved remarkable performance in diverse applications, from code generation (zhang2024codeagent; yang2024swe), web browsing (wei2025browsecomp; yao2022webshop), medical decision-making (heydari2025anatomy; mcduff2025towards; kim2024mdagents), finance (yu2025finmem), sustainability (zhang2025towards), to scientific discovery (gottweis2025towards; mitchener2025kosmos). As tasks grow in complexity and require sustained environmental interaction, the field has increasingly turned to multi-agent systems (MAS), relying on the premise that specialized collaboration consistently outperforms single-agent systems (SAS) (tran2025multiagent; guo2024large). Previous work has made positive claims about multi-agent systems: “More agents is all you need(li2024more), suggesting that agent collaboration follows collaborative scaling principles (qian2024scaling), and that MAS consistently outperforms single-agent systems (SAS) on complex tasks (du2023improving; chen2023agentverse). Yet, despite rapid adoption, there remains no principled quantitative framework to predict when adding agents amplifies performance and when it erodes it. This gap leaves practitioners relying on heuristics, hindering both the emergence of a science of agent systems and, critically for real-world deployment, the ability to determine when multi-agent coordination provides genuine value over simpler single-agent alternatives.
@@
In contrast to prior claims that “more agents is all you need”, our evaluation reveals that the effectiveness of multi-agent systems is governed by quantifiable trade-offs between architectural properties and task characteristics. We establish a predictive framework using empirical coordination metrics: efficiency (success/overhead ratio), error amplification factors, message density and redundancy, achieving cross-validated $R^{2}{=}0.524$ (explaining more than half of the performance variance on held-out data) without dataset-specific parameters. Critically, this framework generalizes beyond training configurations: the model correctly predicts optimal architectures for 87% of held-out task configurations, demonstrating extrapolation to unseen task structures.
@@
Our analysis identifies three patterns. First, a tool-coordination trade-off ($\beta{=}{-}0.267$, p<0.001p{<}0.001): tool-heavy tasks (e.g., 16-tool software engineering) suffer from multi-agent coordination overhead, with efficiency penalties compounding as environmental complexity increases. Second, a capability ceiling ($\beta{=}{-}0.404$, p<0.001p{<}0.001): tasks where single-agent performance already exceeds 45% accuracy experience negative returns from additional agents, as coordination costs exceed diminishing improvement potential. Third, we observe architecture-dependent error amplification. Independent systems amplify errors $17.2\times$ through unchecked error propagation, where individual mistakes cascade to the final output. Centralized coordination, however, contains this to $4.4\times$ by enforcing validation bottlenecks that intercept errors before aggregation. Performance spans $81\%$ relative improvement (structured financial reasoning under centralized coordination) to $70\%$ degradation (sequential planning under independent coordination), demonstrating that architecture-task alignment, not number of agents, determines collaborative success. Importantly, optimal architectures vary systematically: decentralized coordination benefits tasks requiring parallel exploration of high-entropy search spaces (dynamic web navigation: $9.2\%$), while all multi-agent variants universally degrade performance on tasks requiring sequential constraint satisfaction (planning: $39\%$ to $70\%$), where coordination overhead fragments reasoning capacity under fixed computational budgets. We synthesize these findings into quantitative architecture selection rules (Section 4.3) achieving 87% prediction accuracy on held-out configurations. The underlying mechanisms driving these patterns are interpretable: the tool-coordination trade-off arises because multi-agent systems fragment the per-agent token budget, leaving insufficient capacity for complex tool orchestration; the capability ceiling reflects that coordination overhead becomes a net cost when baseline performance is already high; and architecture-dependent error amplification stems from the presence or absence of validation bottlenecks that catch errors before propagation. These mechanistic insights enable practitioners to move from architectural heuristics to principled, measurement-driven deployment decisions.
@@
  • Formalization of Agentic Evaluation rigor: We redefine rigorous agentic assessment by distinguishing it from static reasoning tasks (e.g., MMLU). We establish that valid agentic evaluation requires three necessary conditions: sustained multi-step environment interaction, iterative information gathering under partial observability, and adaptive strategy refinement based on feedback.
@@
  • Intelligence-Coordination alignment: We characterize the non-linear relationship between foundational model capabilities and agentic performance. We demonstrate that while higher capability (Intelligence Index) offers accelerating returns, these gains are not automatic; they strictly depend on architectural alignment. Without correct coordination structures, foundational improvements are often negated by coordination overhead.
  • Quantitative scaling principles and architecture alignment: We derive a mixed-effects model ($R^{2}{=}0.524$) using empirical coordination metrics—efficiency ($E_{c}$), error amplification ($A_{e}$), and redundancy ($\rho$) to quantify how performance emerges from the interplay of reasoning capability and task properties. This framework identifies fundamental limits on coordination, specifically a tool-coordination trade-off ($\beta{=}{-}0.267$) where tool-heavy workflows suffer from coordination tax, and safety bounds where centralized verification reduces error amplification from $17.2\times$ to $4.4\times$. Leveraging these mechanisms, we demonstrate that architecture selection is governed by measurable task features (e.g., decomposability) rather than simple agent scaling, achieving 87% accuracy in predicting optimal architectures on held-out tasks.
@@ 3.1 System Definition
The observation history space $\mathcal{H}$ contains sequences of action-observation pairs. The decision function $\pi_{i}$ is instantiated by the reasoning policy $\Phi_{i}$ (the LLM): given a history $h_{i,t}$, the LLM generates a reasoning trace and selects the next action.
For instance, a history $h_{i,t}=[(\text{``search(query=`pandas')''},\text{``Found 5 files''}),$...] is processed by $\Phi_{i}$ to produce the next tool call $\alpha_{i,t+1}$.
@@
  • Independent: $C=\{(a_{i},a_{\text{agg}}):\forall i\}$ (agent-to-aggregator only, no peer communication)
@@
  • Independent MAS: $A=\{a_{1},\ldots,a_{n}\}$, $\mathcal{C}=\{(a_{i},a_{\text{agg}})\}$, $\Omega=\texttt{synthesis\_only}$. The synthesis_only policy concatenates sub-agent outputs without cross-validation or majority voting; the aggregator performs no analytical comparison of responses, ensuring that any performance differences arise purely from parallel exploration rather than error correction. This achieves maximal parallelization but minimal coordination, suitable for ensemble-style reasoning.
@@ 4 Experiments & Results
Figure 2: Comparative performance of single-agent system (SAS) and multi-agent system (MAS) across four agentic benchmarks reveals highly task-dependent scaling dynamics. Box plots show distribution of success rates (scale: 0 to 1, where 1 represents 100% success). Percentage annotations represent relative improvement/degradation compared to SAS baseline: $(\text{mean}_{\text{MAS}}-\text{mean}_{\text{SAS}})/\text{mean}_{\text{SAS}}\times 100\%$. SAS serves as the reference baseline (shown without percentage annotation). (a) BrowseComp-Plus shows polarized results, with independent agents catastrophically underperforming relative to SAS (-35%) while more structured coordination achieves modest gains. (b) Finance Agent demonstrates the strongest multi-agent benefits, with all MAS architectures substantially outperforming SAS (from +57 to 81%), suggesting that complex planning and distributed reasoning provide significant advantages in structured economic domains. (c) PlanCraft exhibits consistent degradation across all MAS variants (from -70% to -39%). The core difference from Finance Agent lies in task structure, where Finance Agent tasks decompose into parallelizable subtasks (e.g., separate agents can independently analyze revenue trends, cost structures, and market comparisons, then synthesize findings), whereas PlanCraft requires strictly sequential state-dependent reasoning, each crafting action modifies the inventory state that subsequent actions depend upon. (d) Workbench shows marginal effects (from -11 to +6%), suggesting balanced trade-offs between problem structure and orchestration costs. White diamond markers denote per-architecture mean performance.
@@ 4.1 Setup
We evaluate three LLM families across multiple model sizes, spanning externally standardized Intelligence Index values from 42 to 71 (a composite capability score integrating reasoning, coding, and knowledge benchmarks; see Appendix A):
@@
We tested five coordination topologies: Single-Agent System (SAS) and four Multi-Agent System (MAS) variants: Independent, Centralized, Decentralized, and Hybrid. Rather than attempting exhaustive coverage of all possible architectures, we selected these four MAS configurations to form a structured ablation over two key coordination dimensions: (i) orchestrator presence (hierarchical control vs. flat structure), and (ii) peer communication (direct sub-agent interaction vs. isolated execution). Independent isolates pure ensemble effects without any inter-agent communication; Centralized introduces hierarchical verification through an orchestrator bottleneck; Decentralized enables peer-to-peer information fusion without hierarchy; and Hybrid combines both mechanisms (see Table 2 for formal complexity characterization). This design enables causal attribution of performance differences to specific coordination mechanisms rather than generic “multi-agent” effects. Coordination complexity is parameterized by communication overhead: the total number of inter-agent message exchanges required per task, yielding empirical values ranging from 0% (SAS) to 515% (Hybrid), with Independent at 58%, Decentralized at 263%, and Centralized at 285% relative to the single-agent baseline (see Table 5).
@@
Primary outcome is task success/accuracy (domain-dependent: factual correctness for Finance Agent, task completion for Workbench, goal satisfaction for PlanCraft, page synthesis accuracy for BrowseComp-Plus). Secondary metrics include: (i) factual error rate $E$ via domain-specific validators (Cohen’s $\kappa$ (cohen1960coefficient): Finance Agent =0.91=0.91, Workbench =0.89=0.89, PlanCraft =0.87=0.87, BrowseComp-Plus =0.88=0.88; exceeding 0.80, indicating strong inter-rater reliability); (ii) information gain $\Delta\mathcal{I}$ from pre- vs. post-coordination uncertainty proxies (see Eq. 2); (iii) token-overlap structure across agent rationales, labeling tokens as unique (appearing in exactly one agent), shared (two or more agents), or contradictory (semantic opposition detected when BERTScore similarity <0.3<0.3 between assertion pairs, i.e., $1-\text{BERTScore}>0.7$, following the dissimilarity threshold established by zhang2019bertscore); (iv) efficiency metrics including success per 1,000 tokens and cost-normalized performance. All metrics are normalized per reasoning turn and per token to enable cross-architecture comparison. We select coordination metrics based on two criteria: (i) direct measurability from experimental traces without requiring ground-truth labels beyond task success, and (ii) coverage of distinct aspects of coordination–performance relationships identified in prior work (cemri2025why). We excluded metrics requiring subjective human annotation (e.g., solution creativity) or those exhibiting high collinearity with included measures (e.g., total message count correlates $r>0.92$ with overhead). Variance inflation factor (VIF) analysis confirmed no severe multicollinearity among retained predictors (all VIF <5<5). Specifically:
@@ 4.2 Main Results
Multi-agent systems demonstrate highly heterogeneous performance across task domains, contingent on problem structure and architectural choices. On Finance Agent, MAS achieve substantial improvements: Centralized reaches +80.8% (mean 0.631 vs. SAS 0.349), Decentralized achieves +74.5% (0.609), and Hybrid reaches +73.1% (0.604), driven by opportunities for distributed financial reasoning across multiple agents. On Workbench, multi-agent systems show minimal gains: Decentralized achieves +5.7% (0.664 vs. SAS 0.629), while Centralized and Hybrid both slightly underperform at -1.2%. On BrowseComp-Plus, improvements remain modest: Decentralized achieves +9.2% (0.347 vs. SAS 0.318), with Centralized essentially flat at +0.2%. Critically, PlanCraft exhibits universal performance degradation across all multi-agent architectures. Centralized declines to 50.3-50.3% (0.282 vs. SAS 0.568), Decentralized to 41.5-41.5% (0.332), Hybrid to 39.1-39.1% (0.346), and Independent to 70.1-70.1% (0.170). To understand this stark contrast between Finance Agent’s gains and PlanCraft’s degradation, we examined execution traces from both domains. In PlanCraft, efficient single-agent trajectories follow direct execution paths. For example, crafting a diorite_wall:
Turn 1: search("diorite_wall") $\rightarrow$ Recipe: 6 diorite in 2x3
Turn 2: move(diorite $\rightarrow$ crafting_grid)
Turn 3: craft $\rightarrow$ Task complete
In contrast, centralized multi-agent systems decompose inherently sequential tasks into artificial subtasks:
Agent 1: Research recipe (redundant—lookup is instantaneous)
Agent 2: Check inventory (redundant—state visible to all)
Agent 3: Execute crafting (the only necessary step)
This unnecessary decomposition generates substantial coordination messages on average for tasks requiring only a few execution steps, consuming token budget on coordination rather than reasoning. Conversely, Finance Agent trajectories demonstrate when coordination provides genuine value. Single-agent execution exhibits sequential bottlenecks:
Turn 1: web_search("merger news") $\rightarrow$ Surface results
Turn 2: edgar_search("filings") $\rightarrow$ Limited depth
Turn 3--7: Sequential exploration with insufficient breadth
Centralized coordination enables parallel information synthesis:
Agent 1: Regulatory/news analysis
Agent 2: SEC filing research
Agent 3: Operational impact assessment
Orchestrator: Synthesize multi-source findings
The task’s natural decomposability such as revenue, cost, and market factors can be analyzed independently which aligns with the coordination structure, yielding +80.9+80.9% improvement. These trajectory patterns reveal the mechanistic basis for domain-dependence: coordination overhead becomes counterproductive when coordination complexity exceeds task complexity (PlanCraft), but provides substantial gains when tasks naturally decompose into parallel information streams (Finance Agent).
Aggregating across all benchmarks and architectures, the overall mean MAS improvement is 3.5-3.5% (95% CI: [18.6-18.6%, +25.7+25.7%]), reflecting substantial performance heterogeneity with high variance ($\sigma=45.2$%). The performance range across MAS variants spans from 70.0-70.0% (PlanCraft Independent) to +80.9+80.9% (Finance Centralized), indicating that MAS do not provide universal benefits but rather domain-specific trade-offs.
@@
Mixed-effects regression confirms domain complexity (refer to Appendix C for more details) as a significant negative moderator of MAS advantage ($\hat{\beta}=-0.114$, $95\%$ CI: [0.186,0.042][-0.186,-0.042], $p=0.002$). The mechanism operates through fixed computational budgets (matched total tokens across MAS and SAS): in structured, decomposable domains (Finance Agent, moderate Workbench instances), agents complete local reasoning with residual capacity available for inter-agent communication. Here, inter-agent messages reduce variance through redundancy elimination and enable synthesis of partial solutions, producing large performance deltas (Finance: $80.9\%$). Conversely, in high-complexity sequential domains (PlanCraft), intra-agent reasoning for constraint verification and state tracking consumes most available tokens before communication can occur; subsequent inter-agent messages then compress reasoning quality and produce strong negative returns (PlanCraft: $39.0\%$ to $70.0\%$).
@@
While domain complexity broadly moderates MAS effectiveness, the architecture-domain interaction reveals non-uniform preferences even within similar complexity regimes: no single architecture dominates across all domains and vendors. Architecture effectiveness depends critically on domain structure: Finance Agent benefits most from Centralized (+80.9%) and Decentralized (+74.5%), Workbench from MAS-Decentralized (+5.6%), and BrowseComp-Plus from MAS-Decentralized (+9.2%). In degrading domains, architecture selection becomes a least-worst optimization: PlanCraft shows Hybrid as relatively best (-39.0%) compared to MAS-Centralized (-50.4%) and MAS-Independent (-70.0%).
Family-specific coordination preferences emerge within improvement-positive domains. On Finance Agent, Anthropic’s MAS-Centralized achieves +127.5% (0.636 vs. 0.280 SAS), indicating conservative but stable coordination, whereas Google’s MAS-Centralized reaches +164.3% (0.740 vs. 0.280 SAS, averaging Centralized performance), suggesting stronger attention-mechanism alignment with hierarchical message exchange; OpenAI’s MAS-Centralized achieves +69.9% (0.79 vs. 0.465 SAS). On Workbench, where multi-agent overhead is less tolerable (efficiency degrades from $E_{c}=0.466$ for SAS to $E_{c}=0.074$ for Hybrid, the largest relative drop across benchmarks), Anthropic’s best variant (MAS-Decentralized, +10.8%) remains superior to Google (+9.5%) and OpenAI (+8.6%), reflecting relative efficiency in managing coordination costs. Critically, on PlanCraft where all variants degrade, vendor preferences flatten: Anthropic shows maximum -54.5% (MAS-Hybrid 0.31 vs. SAS 0.68), Google shows -25.3% (best), and OpenAI shows -32.3%, indicating that communication mechanisms cannot overcome fundamental sequential reasoning constraints. While the precise mechanisms remain to be characterized, potential factors include differences in instruction-following fidelity, context utilization patterns, and inter-turn consistency that affect how agents interpret and respond to coordination messages. No vendor achieves universal multi-agent dominance; instead, each exhibits relative advantages in structured domains (Finance) that evaporate in sequential constraint-satisfaction domains (PlanCraft), indicating that multi-agent benefits are genuinely contingent on problem structure rather than generalizable across task types.